OpenStudy (anonymous):

solve for x (4/9)^sqrt of x = (2/3)^39 - sqrt of x

OpenStudy (asnaseer):

Please confirm if this is what you want to solve? \[(4/9)^{\sqrt{x}}=(2/3)^{39}-\sqrt{x}\]

OpenStudy (anonymous):

sorry, 39 take the sqrt of x is the power

OpenStudy (king):

Can you sbmit as a drawing

OpenStudy (asnaseer):

Is this correct? \[(4/9)^{\sqrt{x}}=(2/3)^{39-\sqrt{x}}\]

OpenStudy (anonymous):